A coin is tossed twice at random what is the probability of getting

(a)At least one tail

(b)The same face


1
Expert's answer
2021-06-29T07:55:07-0400

Solution:

S={HH,HT,TH,TT}

(a) Event of at least one tail = {HT,TH,TT}

P(at least one tail)"=\dfrac{3}{4}"

(b) Event of the same face = {TT,HH}

P(The same face)"=\dfrac{2}{4}=\dfrac{1}{2}"